### v3

# LiLIS: Enhancing Big Spatial Data Processing with Lightweight Distributed Learned Index

Abstract
The efficient management of big spatial data is crucial for location-based services, particularly in smart cities. However, existing systems such as Simba and Sedona, which incorporate distributed spatial indexing, still incur substantial index construction overheads, rendering them far from optimal for real-time analytics. Recent studies demonstrate that learned indices can achieve high efficiency through a well-designed machine learning model, but how to design learned index for distributed spatial analytics remains unaddressed. In this article, we present LiLIS, a Lightweight distributed Learned Index for big Spatial data. LiLIS amalgamates machine-learned search strategies with spatial-aware partitioning within a distributed framework, and efficiently implements common spatial queries, including point query, range query, $k$-nearest neighbors ($k$NN), and spatial joins. Extensive experimental results over real-world and synthetic datasets show that LiLIS outperforms state-of-the-art big spatial data analytics by 2-3 orders of magnitude for most spatial queries, and the index building achieves 1.5-2$\times$ speed-up. The code is available at https://github.com/SWUFE-DB-Group/learned-index-spark.
Index Terms:
Learned index, spatial data, spatial databases, distributed analytics.
I Introduction
Tthe widespread adoption of GPS-enabled devices has led to the generation of extensive spatial data, enabling the collection and dissemination of trajectory information through location-based services (LBS) such as Google Maps, Uber, and geo-tagged social media platforms [1, 2, 3, 4]. This development provides researchers with unprecedented opportunities to mine large scale spatial data. For example, DiDi Chuxing [^1] https://web.didiglobal.com/, a leading ride-sharing platform in China, receives more than 25 million trip orders and produces over 15 billion location points per day [5]. To efficiently manage the voluminous spatial datasets, technological innovations in terms of distributed storage and processing have been proposed to handle the limitation of a single-node machine [6, 7]. State-of-the-art big spatial data analytics, such as Simba [8] and Sedona [9], are able to provide high throughput by providing inherent spatial indexing and optimization. Nevertheless, current solutions which rely on traditional spatial indices such R-tree [10] and Quadtree [11] are still hindered by considerable index construction overhead and runtime query cost. Therefore, developing efficient approaches to manage and process large-scale spatial data remains a critical challenge for modern data systems [12].
In recent years, learned indices, which utilize machine learning (ML) models to predict the data position directly, have paved the way to efficient search with a succinct data structure [13, 14]. Subsequently, various learned indices over spatial data are proposed to support two-dimensional spatial data access [15, 16, 17, 18, 19], reducing the search time from $O(\log{N})$ to $O(1)$ given $N$ spatial objects. At a high level, a learned index can be seen as a function which is able to map the key (e.g., spatial coordinates in our work) to the position of underlying data directly with a pre-defined error bound, as illustrated in Figure 1. However, to the best of our knowledge, there is no any big spatial system that is integrated with learned index.
Figure 1: An illustration of a learned index in which the key is a spatial coordinate $(x,y)$ with a value $v$, and a model maps the key to the position in memory/disk directly with an error bound.
There are several key challenges to design learned spatial indices for distributed analytics. Firstly, training a machine learning model to achieve high-quality predictions is no trivial task, as it often involves a time-consuming training process, especially when dealing with large-scale datasets (C1). Secondly, most model-based learned index methods are not compatible with existing query paradigms, necessitating significant refactoring efforts (C2). Last but not the least, existing learned index methods support only a limited range of spatial queries, restricting their applicability in diverse real-world scenarios (C3). To bridge the gap mentioned above, we present a Lightweight distributed Learned Index for big Spatial data systems, dubbed as LiLIS. Specifically, In order to address those challenges, several novel techniques in terms of distributed learned spatial index are developed in this paper:
  • We leverage an error-bounded spline interpolation to implement LiLIS by learning the two-dimensional distribution of the underlying spatial data. This learned index can be realized with very few parameters using simple yet effective interpolating techniques, making the spatial index considerable lightweight compared to traditional spatial indices (e.g., R-tree and KD-tree), and it also serves as the foundation for efficient spatial queries. (C1)
  • Unlike other learned indices, LiLIS is compatible with existing execution engines of big data systems, and thus provides native features such as scalability and fault tolerance. For example, LiLIS can utilize spatial-aware partitioning and two-phase pruning to boost the performance. The LiLIS prototype implemented on Apache Spark is straightforward without hacking the underlying execution engine. (C2)
  • LiLIS employs a novel decoupling of data partitioning and indexing components, enabling seamless extension to support diverse spatial query types including point queries, range queries, $k$-nearest neighbors ($k$NN), and spatial joins. Notably, while maintaining the same learned index model architecture, LiLIS is able to adapt different partitioning strategies based on data and query distributions. (C3)
In summary, our contributions are threefold:
  1. 1.
    We designed LiLIS, a lightweight distributed learned index for big spatial data. To the best of our knowledge, this is the first work to introduce spatial learned indices to big data analytics.
  2. 2.
    We implemented efficient distributed spatial algorithms to support common spatial queries, including point query, range query, $k$NN, and spatial join based on LiLIS on Apache Spark.
  3. 3.
    Our comprehensive experimental evaluation on both real-world and synthetic datasets demonstrates LiLIS’s significant performance advantages. The results indicate that LiLIS achieves a 2-3 order of magnitude speed improvement over state-of-the-art big spatial analytics systems for most queries, while also delivering 1.5-2× faster index construction.
The remainder of this article is organized as follows: Section II reviewed the related work. The overview design and key methodology of LiLIS are described in Section III. We discussed the details of efficient distributed spatial algorithms in Section IV. Section V presents the results of our experiments, providing insights about the efficiency and scalability of LiLIS, and Section VI concludes this article and suggests directions for future research.
II Related Work
II-A Spatial Data Indexing
A spatial index is a carefully designed structure engineered to manage spatial objects by leveraging their spatial attributes, such as location, geometry, properties, and relationships [20]. Spatial indices are widely studied and commonly implemented in mainstream spatial databases, with tree-based index structures (e.g., KD‑tree [21], Quadtree [22], and R-tree [23]) being the most commonly used. Tree-based spatial indices are typically designed as hierarchical structures, partitioning space recursively. Consequently, for most queries, they can generally achieve an average-case time complexity of $O(\log{N})$ , where $N$ represents the number of spatial objects. In addition to tree‐based index structures, a simple yet effective technique is grid partitioning, which divides the two‐dimensional data space into fixed‐size or adaptively adjusted grid cells [24, 25].
Conventional spatial indexing schemes mentioned above are typically designed as general-purpose structures without optimizations aligned to particular data distribution patterns or domain-specific characteristics. Consequently, when applied to large-scale, high-dimensional datasets, they frequently suffer from reduced query throughput and incur significant storage and I/O overheads [26, 27, 28]. Therefore, there is a compelling need for a lightweight yet efficient solution to mitigate these challenges.
II-B Learned Indices
A learned index is an advanced methodology that leverages machine learning models to replace conventional database indexing structures, such as B-trees and hash tables. The core principle of a learned index is to develop a function that maps a search key to the corresponding storage address of a data object, thereby optimizing query performance [14, 29, 30, 31]. T. Kraska et al. [13] first introduced the concept of learned indexes, and they proposed the Recursive Model Index (RMI), which hierarchically models data distributions through a multi-stage neural network [32]. However, while their foundational work demonstrates that a learned index can deliver up to a threefold reduction in search time and an order-of-magnitude decrease in memory footprint [13], it remains confined to static, read-only workloads. Subsequently, J. Ding et al. introduced ALEX [33], an updatable learned index that adaptively partitions data to sustain high query and insert performance under dynamic workloads. Beyond the one-dimensional learned indices, later work further extends them into multi-dimensional settings [34]. The Z-order model [16] extends RMI to spatial data. Their method uses a space-filling curve to order points and learns the Cumulative Distribution Function (CDF) to map each point’s key to its rank. The recursive spatial model index [35] further develops these principles by employing the rank space–based transformation technique [36].
Although research on learned indices has significantly enhanced query performance, current studies are confined to single-machine environments and fall short in addressing large-scale data processing. At present, there is a significant research gap in the field of distributed learned indices.
II-C Distributed Spatial Analytics
To process large scale spatial data, a large body of big spatial analytics have been proposed over the years [37, 38]. Simba [8] enhances SparkSQL with DataFrame based spatial queries using two-stage RDD indexing. SpatialSpark [39] implements node-local R-trees on Spark RDDs to minimize data transfer. GeoMesa [40] provides spatio-temporal indexing across multiple database backends with Kafka integration. Among these, Apache Sedona [41] stands as state-of-the-art solution, and it delivers superior performance for large-scale geospatial processing through its innovative distributed indexing and query optimization architecture.
Existing distributed spatial systems typically rely on traditional indexing methods [19], such as R-trees or grids, which impose considerable I/O and memory overhead during both construction and querying. However, there is no any distributed spatial analytics equipped with learned spatial indies yet.
III LiLIS Design
Figure 2: The architecture of LiLIS on Apache Spark. It introduces spatial RDD with learned index, and supports a wide range of spatial queries.
In this paper, we mainly adopt Apache Spark, one of the most popular distributed computing engine nowadays, as the underlying framework to implement LiLIS to support spatial points, and the architecture is illustrated in Figure 2. As we can see, it introduces an extra wrapper beyond the resilient distributed dataset (RDD) so that LiLIS is able to support 1) spatial functions (e.g., geometry transformation and distance computation); and 2) spatial predicates (e.g., contains, intersects and within). Notably, a leaned index component is implemented within partitions of spatial RDDs, which are generated by several spatial-aware partition strategies. By decoupling indexing and partitioning, LiLIS applies learned models to replace traditional spatial search with a learned alternative while maintaining compatibility with existing components in distributed frameworks, enabling the feasible implementation of a diverse range of spatial query algorithms. It is worthy to note that the overall design of LiLIS is independent with the specific distributed framework, and the porting to alternatives such as Apache Flink is left as a future work. In what follows, the design of spatial-aware partitioner and learned index are detailed.
III-A Spatial-aware Partitioner
Typically, data in a distributed system shall be partitioned among machines in a cluster. For the sake of load balance and data locality, data-aware partitioners beyond plain range and hash strategies are designed to boost the performance [42]. Therefore, LiLIS adopts the similar method used in Simba [8] to leverage the spatial locality. For the large scale spatial data, we observed that partitioning based on sampling can achieve a great trade-off between efficiency and effectiveness. In this paper, we set sampling rate to 1% in a uniform way. It is worth to note that partitioner itself can be seen as a global index, and LiLIS provides several built-in grid-based (e.g., fixed and adaptive grid) and tree-based (e.g., R-tree, Quadtree, and KD-tree) spatial-aware partitioning methods to adapt for a wide range of applications. As for tree-based partitioners, LiLIS only maintains the leaves as the partitions. When there is no ambiguity, the leaves generated from tree-based methods are also called grids in this paper. Notably, since R-tree is generally built in a bottom-up way, the sampling-based partitioning may not cover all spatial objects. Therefore, we introduce a novel concept in LiLIS, dubbed as an overflow grid, referring to a special grid for all remaining overflowed spatial objects that do not belong to any leaf in R-tree.
Subsequently, each grid is assigned to a unique identifier (an int in this paper) for re-partition. Without loss of generality, we assume the partitioner is R-tree based, and the partitioning algorithm is illustrated in Algorithm 1.
Data: Spatial RDD $R$
1 $SR\leftarrow$ sampling over $R$;
2 Construct grids list $G$ through the specified method over $SR$;
 /* Outer foreach is a parallel map in Apache Spark */
3 foreach object $o\in R$ do
4       isOverflow $\leftarrow True$;
5       foreach grid $(id,g)\in enumerate(G)$  do
6             if $g$ contains $o$ then
7                   $R^{*}$.add($id$, $o$);
8                   isOverflow $\leftarrow False$;
9                   break;
10
11             end if
12
13       end foreach
14      if isOverflow then
15             $R^{*}$.add(len($G$), $o$);
16
17       end if
18
19 end foreach
20Re-partition over $R^{*}$ by keys;
21 Map $(id,o)$ to $o$ over $R^{*}$;
Algorithm 1 Partitioning algorithm in LiLIS
The outer for-loop (Lines 3-15) is indeed a parallel map operation in Apache Spark, and each spatial object $o$ will be mapped into a tuple $(id,o)$, where $id$ is the identifier of the grid in which $o$ locates. The enumerate function borrows from Python (Line 5), emitting a tuple containing an $id$ (starting from 0) and the $grid$ obtained from iterating over $G$. As for overflowed object, its partition $id$ will be assigned to the size of grids (Lines 12-14). Therefore, a spatial RDD $R$ will be transformed into a key-value RDD $R^{*}$, and a re-partitioning procedure (Line 16) through shuffling can be conducted over $R^{*}$ by the key (i.e., the $id$ of the partition). Finally, $R^{*}$ will be mapped back into a spatial RDD with objects only (Line 17). Since the partitioner acts as a global index, the master node must maintain all partitions’ properties within $G$, including their spatial boundaries and identifiers.
III-B Distributed Learned Spatial Index
To keep the compatibility with existing systems [8, 9, 42], LiLIS follows the two-phase filtering solution, and the local index is implemented via a learned model within a given partition. Assume that a spatial object is $(x,y,v)$, where $x$, $y$, and $v$ is its latitude, longitude, and extra attribute (e.g., the textural description), respectively. The two-dimensional spatial points shall be projected into one dimension for sorting, and possible sorting criteria can be either one arbitrary axis or some aggregated value (e.g., Z-order curve and GeoHash). In this paper, we denote the chosen sorting criteria as the key of a spatial object, so each spatial point can be represented as a tuple of $(key,x,y,v)$.
The main idea of the learned spatial index in LiLIS is to construct a spline model so that it can predicate the position of every spatial point with error guarantees [43]. To be specific, given an object $(key,x,y,v)$ whose position in sorted dataset $\mathcal{D}$ is $p$, the spline model $\mathcal{S}$ can be described as $\mathcal{S}(key)=\hat{p}\pm\epsilon$, where $\epsilon$ is a pre-specified position error bound so that $|\hat{p}-p|\leq\epsilon$. To build model $\mathcal{S}$ over sorted dataset $\mathcal{D}$, we need to find a set of representative objects so that the maximum interpolation error is not larger than $\epsilon$, as illustrated in Figure 3. As for the one-pass error-bounded spline algorithms that construct the model, we refer the readers to [44]. Given a partition $P$ with $N$ spatial objects, the distributed learned spatial index is conducted by mapPartitions in Apache Spark, without time-consuming data shuffling between machines.
Figure 3: The main idea of spline index is to obtain an estimated position $\hat{p}$ of a given key $k$ based on two adjacent points (Step 1), and then perform a binary search within positions $\hat{p}\pm\epsilon$ in sorted dataset (Step 2). The overall searching time complexity is constant after retrieving the lower bound $k_{l}$ and upper bound $k_{r}$.
Compared to the commonly-used R-tree index using Sort-Tile-Recursive (STR) algorithm [45] whose time complexity is $O(N\log{N}+N\log{f}\times\log_{f}{N})$, where $f$ is the fanout, the time complexity of the learned index in LiLIS is $O(N\log{N}+N)$. As for searching, the time complexity of the learned index is $O(\log{M})$, where $M\ll N$ is the size of spline points set. As an optimization, the searching time can be further deduced to be constant on average if a radix table is introduced [43].
Input: Spline points $S$, number of radix bits $b$
Output: Radix table $T$
1 $T\leftarrow$ new int[1 $\ll$ b + 2];
2 $min,max\leftarrow S.first.key,S.last.key$;
3 $f\leftarrow(1\ll b)/(max-min)$;
4 $T[0]\leftarrow 0$;
5 $prev\leftarrow 0$;
6 foreach point $(i,p)\in enumerate(S)$ do
7       $curr\leftarrow(int)((p.key-min)\times f)$;
8       foreach $j\in[prev+1,curr]$ do
9             $T[j]=i$;
10
11       end foreach
12
13 end foreach
[…]
15       $T[prev+1]\leftarrow n-1$;
16
17 }}\}}
18return T;
Algorithm 2 Building a radix table for floating keys
The radix table was proposed to handle the unsigned integer keys (e.g., uint32 and uint64) only. To this end, LiLIS extends this design to support floating numbers and strings to fit spatial data. Algorithm 2 outlines how to build a radix table for floating keys. The size of the radix table is bounded by a user-defined number of bits $b$. Generally, this radix table is to compress the spline points set (Lines 6-14) with a scaling factor $f$ (Line 3). Therefore, given a search key $k$, let $k^{\prime}$ be $(int)(k-min)\times f$, the lower bound and upper bound position are $S[T[k^{\prime}]]$ and $S[T[k^{\prime}+1]]$, respectively.
Remark 1: Converting a string to an unsigned integer is a well-established process, as it can be efficiently accomplished by leveraging robust, thoroughly validated hash code computation techniques, such as the product-sum method, which is natively implemented in programming languages like Java.
III-C Discussions on LiLIS Design
III-C1 Enhancing Rather Than Replacing
Prior work on learned indices often focuses on replacing traditional indexing structures, which necessitates significant refactoring of system architectures and algorithm implementations. In contrast, LiLIS adopts a complementary approach by enhancing the existing computation engine without requiring disruptive overhauls, thus reducing adoption barriers for real-world systems.
III-C2 Lightweight
Motivated by the objective of enhancing efficiency, LiLIS prioritizes a lightweight design. Consequently, constructing a learned index with LiLIS entails minimal overhead due to its reliance on a limited number of parameters. Furthermore, an additional benefit lies in the ability to develop a dual-indexing framework, enabling the selection of an appropriate index based on specific requirements. For instance, when either $x$ or $y$ serves as the key, the construction of an R-tree and a learned index can leverage a shared sorting procedure, thereby optimizing the indexing process.
IV Query Algorithms in LiLIS
In this section, we discuss how to design efficient query algorithms based on the learned index proposed in Section III.
IV-A Point Query
Given a query point $q=(x,y)$ and dataset $\mathcal{D}$, a point query returns true if $q$ is within $\mathcal{D}$, and returns false otherwise. Obviously, there is at most one partition that is overlapped with $q$, and this can be done with RDD’s filter method. If the candidate partition exists, then we use the learned index to predict the estimated position, and refine the final result according to Algorithm 3. As illustrated in Figure 3, we first compute the estimated position (Line 1), and then search the key within the error bound (Line 2). If there is no such key, the algorithm returns false directly (Lines 3-5). Otherwise, we scan in both directions (Lines 6-17).
Input: Candidate partition $\mathcal{P}$, and query point $q$
Output: True if $q\in\mathcal{P}$, Flase otherwise
1 $\hat{p}\leftarrow\mathcal{P}.learnedSearch(q.key)$;
2 $p\leftarrow$ search $q.key$ over $\mathcal{P}[\hat{p}-\epsilon..\hat{p}+\epsilon]$;
3 if $p==null$ then
4       return False;
5
6 end if
[…]
8 while $p<\mathcal{P}.size$ and $\mathcal{P}[p].key==q.key$ do
9       if $q.x==\mathcal{P}[p].x$ and $q.y==\mathcal{P}[p].y$ then
10             return True;
11
12       end if
13      $p$++;
14
15 end while
[…]
17 while $p\geq 0$ and $\mathcal{P}[p].key==q.key$ do
18       if $q.x==\mathcal{P}[p].x$ and $q.y==\mathcal{P}[p].y$ then
19             return True;
20
21       end if
22      $p$–;
23
24 end while
25return False;
Algorithm 3 Search in point query
IV-B Range Query
Given a rectangle range $q=(x_{l},y_{l},x_{h},y_{h})$, where $(x_{l},y_{l})$ and $(x_{h},y_{h})$ are the lower left and upper right point, respectively, a spatial range query returns all points in $\mathcal{D}$ contained in $q$. Similar to the point query, the globally filtering can be conducted through a linear scan directly. After obtaining the candidate partitions, we apply mapPartitions to compute the resulting points. The overall procedure is similar to Algorithm 3. To be specific, we search from lower bound $(x_{l},y_{l})$ to upper bound $(x_{h},y_{h})$ in a given partition. In practice, it can be optimized from several aspects. For example, if the partition is enveloped entirely within $q$, then all points should be returned without further checking.
Remark 2: The range query above is indeed a rectangle range query. As for a circle range query with a center point $p=(x,y)$ and a radius $r$, a practical solution is to construct a minimal bounding rectangle (MBR) for this circle, and perform a regular rectangle range query first. Finally, a subsequent refinement step is necessary to filter out false positives.
IV-C $k$NN Query
Formally, given a query point $q=(x,y)$, dataset $\mathcal{D}$, and the distance function $d(\cdot,\cdot)$, a $k$NN query returns $\mathcal{R}\subset\mathcal{D}$ so that $|\mathcal{R}|=k$, and for any point $o\in\mathcal{D}\setminus\mathcal{R}$, and point $p\in\mathcal{R}$, it holds that $d(q,p)\leq d(q,o)$. In LiLIS, the distributed $k$NN query is implemented with the range query. Firstly, an estimated range is computed according to the following equations [46]:
$r=\sqrt{k/(\pi d)}$ (1)
where $d$ is density of the dataset, given by
$d=N/\texttt{area}$ (2)
where $N$ and area are the size and area of $\mathcal{D}$, respectively.
The first phase of $k$NN search is to conduct a range query with $\hat{q}=(x_{q}-r,y_{q}-r,x_{q}+r,y_{q}+r)$. If the size of the results is less than $k$, a new searching window is constructed. For $k>1$, the times of calling the range query is theoretically bounded:
$\lceil\frac{\ln{\sqrt{(x_{u}-x_{l})^{2}+(y_{u}-y_{l})^{2}}}-\ln{\sqrt{\frac{k(x_{u}-x_{l})(y_{u}-y_{l})}{\pi N}}}}{\ln{\frac{4k}{\pi(k-1)}}}\rceil$ (3)
where $(x_{l},y_{l})$ and $(x_{u},y_{u})$ are the lower bound and upper bound of $\mathcal{D}$, respectively. For the detailed proof, we refer the reader to [46]. In practice, we found that most $k$NN ($k<10)$ can be answered using no more than twice of range queries.
IV-D Join Query
A spatial join takes two datasets ($\mathcal{A}$ and $\mathcal{B}$) and a spatial predicate ($\theta$) as the input, and it returns a set of pairs in which the predicate holds, i.e., $\sigma_{\theta}(\mathcal{A}\times\mathcal{B})$. Due to the diversity of $\theta$, the implementations of join queries exhibit significant variation. In LiLIS, we currently provide the support for the join between polygons $\mathcal{PG}$ and data points $\mathcal{D}$, and the predicate contains is to check whether a data point is contained in a polygon. For example, the data points can be shops in a city, and the polygons delineate the commercial zones that are selected on-the-fly. A meaningful join query in this context could aim to identify which shops fall within each commercial zone, enabling an analysis of retail density and distribution patterns. In a practical setting, $|PG|\ll|D|$, so polygons can be broadcasted to each partition before performing join. As for the filtering stage, we can compute MBR for a polygon, and then perform a range query whose range is the MBR above to get candiate tuples.
Remark 3: LiLIS also support a few other types of joins (e.g., $k$NN-join) out of the box, but how to leverage the learned index to design a unified framework for more types of joins is left as a future work.
V Experiment
In this section, we conduct extensive experiments to demonstrate the superiority of LiLIS over both real-world and synthetic datasets. All algorithms in this paper were implemented in Java, and all experiments were performed in a cluster with 7 CentOS machines on Tencent Cloud, in which each machine is shipped with 8 GB memory, 80 GB SSD and Intel Duo Processor of 3.0 GHz. As mentioned before, the design of LiLIS is independent with the specific distributed engine, and we present it on Apache Spark 3.0 using Standalone mode, in which one machine is chosen as the master node, and the remaining 6 machines are slave nodes. All configurations in Apache Spark, including garbage collection (GC) and shuffle operations, were maintained at their default settings for the experiments.
V-A Setup
V-A1 Datasets
The datasets used in the experiments are summarized in Table I. Both CHI [^2] https://data.cityofchicago.org/ and NYC [^3] https://www.nyc.gov/site/tlc/about/data.page contain extra information in addition to spatial locations, while SYN consists of spatial points only, which are generated by Spider [^4] https://spider.cs.ucr.edu/. As we can see, although it is possible to load CHI and SYN in a single-node machine, the performance in a distributed setting among different frameworks is a meaningful research question, and the evaluation is enough to demonstrate the effectiveness and efficiency of LiLIS. As for spatial join queries, we constructed 7000 polygons in Chicago, and perform a spatial join with CHI. For other queries, we use NYC as the default dataset.
TABLE I: Spatial Datasets
Name Size #items Description
CHI 1.9 GB 7M Crime events in Chicago
NYC 20 GB 300M New York taxi rides
SYN 3 GB 100M Randomly generated points
Remark 4: While it is true that the default dataset (i.e., NYC) can be accommodated within the memory of a single contemporary machine, the primary objective of this study is to demonstrate the feasibility and efficacy of LiLIS within distributed cloud environments. In such settings, cost-effective commodity hardware with limited memory capacities (e.g., 8 GB), as utilized in our experiments, is typically employed, highlighting the system’s scalability and performance under resource-constrained conditions.
V-A2 Baselines and Implementations
For evaluation, we compare LiLIS with the state-of-the-art big spatial analytics Sedona [9]. Sedona is shipped with two indices (including R-tree and Quadtree) and two partitioners (including KD-tree and Quadtree), and we use Sedona-{I}{P} to denote its various implementation variants, in which {I} and {P} mean the index type and partitioner, respectively. For example, Sedona-RK uses R-tree as the index while adopts KD-tree as the partitioner. The four variants of Sedona are summarized in Table II. Notably, Sedona also allows the no-indexing solution, and we donate it as Sedona-N. As for the implementation using vanilla Apache Spark, we donate it as Spark. As discussed in Section III, LiLIS can adopt different partitioners, we use LiLIS-F, LiLIS-A, LiLIS-Q, LiLIS-K, and LiLIS-R to denote the implementations with fixed-grid, adaptive-grid, quadtree, KD-tree, and R-tree, respectively. In our experiments, we adopt the KD-tree as the default partitioning method for LiLIS. Consequently, LiLIS-K serves as the default implementation in our proposed framework.
TABLE II: Sedona Variants
Index Partitioner KD-tree Quadtree
R-tree Sedona-RK Sedona-RQ
Quadtree Sedona-QK Sedona-QQ
Unlike other learned indices requiring complicated parameters and tuning skills, there are only two hyperparameters in LiLIS, i.e., error bound (default is 32) and number of spline bits (default is 10). The similar setting is also adopted in [43].
V-A3 Workloads
Unlike the workload in Sedona [9] which only focuses on different types of spatial queries, in this paper we further consider more subtle factors for range queries [27]. To be specific, the selectivity, which is the ratio between the area of a query window and the area of the whole dataset, ranges from 0.00001% (default) to 0.1%. In addition, we further adjust skewness for range queries, in which a skewed query (default) follows the distribution of the underlying spatial data, while a uniform query is to construct the input randomly. For $k$NN queries, the default $k$ is 10, ranging from 1 to 1000. To mitigate randomness, all experimental results are averaged over 50 independent runs.
V-A4 Research Questions
In this paper, we mainly focus on the following research questions in terms of LiLIS through comprehensive experimental evaluations:
  • RQ1: How good is LiLIS for common distributed spatial queries compared to the state-of-the-art big spatial analytics?
  • RQ2: What is the effect of different partitioners in LiLIS?
  • RQ3: How good is LiLIS across different datasets?
  • RQ4: What is the effect of subtle factors for range (e.g., varying selectivities) and $k$NN (e.g., varying $k$) queries in LiLIS?
  • RQ5: What is the cost of building indices in LiLIS?
V-B RQ1: Overall Performance
To answer RQ1, we compare LiLIS-K with all other alternative big spatial analytics for range queries, point queries, and $k$NN under the default settings. Because Quad-tree index in Sedona does not support $k$NN, we only report Sedona-RQ and Sedona-RK for $k$NN queries. As for join queries, since the spatial join predicate used in LiLIS is not found in Sedona, we only compare LiLIS-K with Apache Spark. The evaluation results are illustrated in Figure 4. As we can see, LiLIS performs consistently the best compared with other implementations. Since the y-axis in Figure 4 uses a logarithmic scale, the actual performance improvement is significantly greater than the observed values suggest. For example, LiLIS-K takes about 472 millisecond for range queries, while Sedona-RQ takes 521282 millisecond. Interestingly, we can also observe that Sedona variants with indices are not better than Sedona-N and Spark under our settings, and this phenomenon is caused by the underlying data distribution as well as its Filter and Refine model. The authors of Sedona [9] also verify this finding.
Figure 4: The overall performance under default settings.
Takeaway 1 LiLIS is the fastest for spatial points, and it generally outperforms competitors by 2–3 orders of magnitude.
As for RQ1, We also conducted the experiments in terms of throughput (jobs per minute) of LiLIS and its competitors of point and range queries. To be specific, we followed the similar settings in Simba [8], and issued 100 queries on the driver program using a thread-pool with 8 threads. The results are shown in Figure 5. It can be observed that LiLIS-K consistently achieves higher throughput, but the advantage is not as obvious as that in query time, and this is mainly because the throughput is influenced by additional factors such as I/O overhead, CPU cores, and memory bandwidth that are not solely dependent on query execution efficiency.
Figure 5: The throughput (jobs per minute) of point and range queries under default settings.
V-C RQ2: Varying Partitioners
An appropriate partitioner is essential for the best performance. Since the name of a LiLIS variant can indicate the chosen paritioner, we denote the paritioner as its implementation directly in Table III for the sake of simplicity. We highlight the best results in light green and the second-best results in yellow. Our analysis reveals that both Quad-tree and KD-tree serve as highly effective partitioning strategies across a broad range of query types, while the R-tree partitioner emerges as the optimal choice for most queries, with the notable exception of join queries. This distinction underscores a critical tradeoff in spatial indexing: the R-tree’s optimization for efficient individual object retrieval is less effective for join operations, where its structural characteristics are less suited to the combinatorial demands of such queries. In contrast, the more uniform spatial division strategies employed by Quad-tree and KD-tree enable them to handle join queries more adeptly, highlighting their robustness in scenarios requiring complex relational operations.
TABLE III: Varying Partitioners (ms)
Method Query Point Query Range Query $k$NN Query Join Query
LiLIS-F 218.08 704.15 1107.20 636710
LiLIS-A 199.09 521.04 1767.00 1322970
LiLIS-Q 141.3 340.57 773.50 276557
LiLIS-K 82.59 468.64 650.20 228581
LiLIS-R 76.68 471.55 618.80 21492013
Takeaway 2 LiLIS is sensitive to various partitioners, and tree-based partitioners performs better than grid-based partitioners generally.
V-D RQ3: Varying Datasets
As illustrated in Figure 6, the experimental results demonstrate significant performance variations across three query types when executed on three distinct datasets. Point queries exhibited substantial performance disparities, Interestingly, despite the NYC dataset being approximately 10 times larger than CHI, it demonstrated significantly better performance for all queries. We can also notice that queries over SYN consistently performs the best, and this is because uniformly generate spatial objects can be best modeled by both spatial partitioner and learned index. These findings highlight the importance of considering both dataset size and intrinsic characteristics when optimizing spatial database operations. The results demonstrate that larger datasets do not necessarily lead to proportionally longer query times.
Figure 6: LiLIS performance while varying datasets.
The results in Figure 6 also indicate that $k$NN is more time-consuming than both point and range query. Therefore, we further take it as the running example to compare LiLIS with its competitors among various datasets. In Table IV, we highlight the best results in bold, demonstrating that LiLIS consistently outperforms other methods across all datasets.
TABLE IV: Varying Datasets for $kNN$ Queries (ms)
Method Dataset CHI NYC SYN
LiLIS-K 743 650 185
Sedona-RK 7862 790993 83170
Sedona-N 6881 314243 49590
Takeaway 3 LiLIS is affected by data characteristics as well as the size of dataset; LiLIS consistently outperforms other competitors across all datasets.
V-E RQ4: Subtle Factors for Range and $k$NN Queries
As described in the experimental setup, both selectivity and skewness are essential for range queries. To answer those subtle factors, we vary selectivity from 0.000001% to 0.1% under both skewed and uniform queries, and the results are reported in Figure 7. Uniform queries consistently outperform skewed queries due to their alignment with the underlying data distribution, allowing the learned index to make more accurate predictions. Skewed queries exhibit an inverse relationship between selectivity and execution time. To be specific, higher selectivity often reduces query latency. This occurs because skewed queries with small selectivity values frequently target sparse or irregular data regions, leading to inefficient index traversal. In contrast, as selectivity increases, the query workload becomes more balanced, mitigating the impact of skewness.
Figure 7: Skewed and uniform range queries under different selectivities.
Next, we further conduct experiments over all variants of LiLIS while varying $k$ from 1 to 100 for $k$NN queries. According to the results in Figure 8, we can conclude that $k$NN queries is relatively stable when varying $k$, and this is consistent with the analysis presented in Section IV-C. This stability can be attributed to the query processing mechanism: regardless of how $k$ changes, the number of partitions that need to be accessed remains relatively small, as the query radius determination primarily depends on data distribution rather than the $k$ value itself. Additionally, we can also notice that R-tree partitioner consistently outperforms other partitioning strategies across different $k$ values. This superior performance can be explained by the R-tree’s effective spatial indexing capabilities, which enable more efficient pruning of irrelevant partitions during query processing, thereby reducing the overall computational cost.
Figure 8: $k$NN queries in LiLIS when varying $k$.
Takeaway 4 Range queries in LiLIS are sensitive to the skewness, and uniform range queries run stably under different selectivities; $k$NN queries in LiLIS are insensitive to common $k$ ($k<100$).
V-F RQ5: Index Cost
Figure 9: Index building cost.
To demonstrate the superior efficiency of LiLIS in index construction, we conduct comparative experiments measuring the build time for all index variants in both LiLIS and Sedona. As illustrated in Figure 9, LiLIS consistently achieves faster index construction times compared to traditional spatial indices. For instance, building a LiLIS-K index requires approximately 50 seconds, whereas constructing a Sedona-RQ index takes about 100 seconds, representing a 1.5-2.0$\times$ speed-up. Notably, while the performance advantage is significant, the gap in construction time is less pronounced than the query time improvements we observed above. This observation aligns with our theoretical analysis presented in Section III-B.
Takeaway 5 Building indices in LiLIS is faster than in competitors, though the speed-up is less significant than its query performance advantage.
VI Conclusion
Prior work shows that learned indices enable efficient search by using ML models to predict data positions. However, distributed spatial analytics like Sedona still face high index construction overheads, hindering real-time processing. In this study, we propose LiLIS, a lightweight distributed learned index, and develop efficient spatial algorithms to support common spatial queries on Apache Spark, including point queries, range queries, $k$-nearest neighbor ($k$NN) queries, and spatial joins. LiLIS achieves a 2–3 orders of magnitude speed-up over state-of-the-art systems and 1.5-2$\times$ faster index construction. Most notably, LiLIS is the first study to our knowledge to investigate the effectiveness and efficiency of learned index for big spatial data. Future work should investigate alternative indexing approaches (e.g., piece-wise linear model, simple non-linear model and neural network) beyond linear interpolation to make trade-offs between model complexity, accuracy, and inference speed for distributed spatial analytics.